OpenStudy (anonymous):

What is the solution to the equation 36^3c=square root 6^ (8c-6)

OpenStudy (anonymous):

\[36^{3c}+ \sqrt{6}^{8c-6} = 6^{6c} + 6^{4c-3} => 6c=4c-3 => 6c-4c=-3 => c= -3/2\]

OpenStudy (anonymous):

c= -3/2